On a world globe, the distance between Rio de Janerio and Hong Kong, two cities that are actually 17615 kilometers apart, is 21.5 inches. The actual distance between Paris and Stockholm is 1605 kilometers. How far apart are Paris and Stockholm on the same world globe?
Answer by rfer(16322)

17615/21.5=1605/x
17615x=34507.5
x=34507.5/17615
x=1.96 in